OpenStudy (anonymous):

Bob and Susie wash cars for extra money over the summer. Bob's income is determined by f(x) = 6x + 13, where x is the number of hours. Susie's income is g(x) = 4x + 18. If Bob and Susie were to combine their efforts, their income would be h(x) = f(x) + g(x) and would be split evenly between the two of them. Assume Bob works 3 hours. Create the function h(x), and indicate if Bob will make more money working alone or by teaming with Susie.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

How much does Bob earn if he works for 3 hours and he's working alone?

OpenStudy (anonymous):

I'm not sure...

jimthompson5910 (jim_thompson5910):

well x = 3 and f(x) is Bob's income function

OpenStudy (anonymous):

Would I plug that into x?

jimthompson5910 (jim_thompson5910):

so f(3) represents how much money he earns alone

jimthompson5910 (jim_thompson5910):

f(3) = ???

OpenStudy (anonymous):

6x + 13?

jimthompson5910 (jim_thompson5910):

yes, f(x) = 6x + 13

jimthompson5910 (jim_thompson5910):

f(x) = 6x + 13 f(3) = 6(3) + 13 ... replace EVERY copy of x with 3 (since x = 3) f(3) = ??

OpenStudy (anonymous):

So 6(3)+13=31

jimthompson5910 (jim_thompson5910):

he earns $31 on his own

jimthompson5910 (jim_thompson5910):

now calculate h(3) to figure out how much both of them earn together (before they split it up)

OpenStudy (anonymous):

How do I do that?

jimthompson5910 (jim_thompson5910):

h(x) = f(x) + g(x) h(3) = f(3) + g(3) h(3) = ??

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

so you'll need: f(3) and g(3) you just found f(3). What is g(3) ?

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

4(3)+18=30

jimthompson5910 (jim_thompson5910):

so h(3) = ??

OpenStudy (anonymous):

30(3)+31(3)?

jimthompson5910 (jim_thompson5910):

remember we had h(3) = f(3) + g(3)

jimthompson5910 (jim_thompson5910):

and we found f(3) = 31 and g(3) = 30

OpenStudy (anonymous):

So 61?

jimthompson5910 (jim_thompson5910):

since f(3) = 31, you replace all of "f(3)" with "31"

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

that's their total combined income, but now we have to split it evenly amongst 2 people

OpenStudy (anonymous):

How would I do that?

jimthompson5910 (jim_thompson5910):

splitting is the same as dividing

OpenStudy (anonymous):

Right, I got 15.5. But what do I do next, I mean?

jimthompson5910 (jim_thompson5910):

example: you have 10 dollars to split amongst 2 people, so each gets 10/2 = 5 dollars

jimthompson5910 (jim_thompson5910):

15.5 is too low

OpenStudy (anonymous):

Oh sorry!

jimthompson5910 (jim_thompson5910):

you take the combined income and divide it by 2

OpenStudy (anonymous):

30.5

jimthompson5910 (jim_thompson5910):

so which is better? Bob working on his own? or working with Susie?

OpenStudy (anonymous):

On his own?

jimthompson5910 (jim_thompson5910):

why on his own

OpenStudy (anonymous):

Because he would receive more?

jimthompson5910 (jim_thompson5910):

how much more

OpenStudy (anonymous):

One dollar?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

I'm so sorry , this is very new to me.

jimthompson5910 (jim_thompson5910):

what is f(x) again?

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

what does the function represent? Whose income?

OpenStudy (anonymous):

Bob's

jimthompson5910 (jim_thompson5910):

so f(3) represents what?

OpenStudy (anonymous):

The amount of hours he worked?

jimthompson5910 (jim_thompson5910):

x = 3 is the number of hours he worked (3 hours)

jimthompson5910 (jim_thompson5910):

f(3) is the income he gets for working 3 hours

OpenStudy (anonymous):

Right!

jimthompson5910 (jim_thompson5910):

which was what?

OpenStudy (anonymous):

31

jimthompson5910 (jim_thompson5910):

so comparing 31 to 30.5, he loses $0.50 or 50 cents It turns out that as x gets bigger, his loss gets bigger as well. So he's better off working alone.

OpenStudy (anonymous):

Okay, got it. What would the equation be?

OpenStudy (anonymous):

It's either h(x)=2x+5 or h(x)=10x+31

jimthompson5910 (jim_thompson5910):

well remember that h(x) = f(x) + g(x)

jimthompson5910 (jim_thompson5910):

you can replace f(x) with 6x+13 and replace g(x) with 4x+18 then simplify

OpenStudy (anonymous):

(6x+13)+(4x+18)=10x+31

OpenStudy (anonymous):

Jim thank you so much for your patience! I'm so grateful for you guys on OpenStudy!

jimthompson5910 (jim_thompson5910):

you're welcome